kingofdumb:

If the quantity 4 times x times y cubed plus 8 times x squared times y to the fifth power all over 2 times x times y squared is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of a?

Unnecessary:

Any ideas?

AZ:

We want to simplify the left side so that way we can find out what the exponents on 'x' and 'y' are on the right side

AZ:

\(\Large\dfrac{4xy^3 + 8x^2y^5}{2xy^2} = 2x^{a}y^{b} + 4x^{c}y^{d}\)

AZ:

There's a few rules you need to know \( \dfrac{a+b}{c} = \dfrac{a}{c} + \dfrac{b}{c}\)

AZ:

so knowing that, we can re-write \(\large\dfrac{4xy^3 + 8x^2y^5}{2xy^2} \) as \(\large\dfrac{4xy^3}{2xy^2} + \dfrac{8x^2y^5}{2xy^2} \) and so now here's an exponents rule that you should know \(\large \dfrac{a^b}{a^c} = a^{b-c}\) Can you simplify it further? how would you divide \(\dfrac{4xy^3}{2xy^2}\) just do it part by part what is 4/2 = ? what is x/x = ? what is \(\dfrac{y^3}{y^2} = ?\)